OpenStudy (anonymous):

Please help I'm so freakin confused!!!!! Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses.

OpenStudy (anonymous):

\[f(x)=\frac{ x+a }{ b }\] and \[g(x)=cx-d\]